OpenStudy (anonymous):

Find the x-coordinates of the points of intersection of the graphs of the given equations. y^2-2x^2=3 x+y=3

OpenStudy (anonymous):

\[y^2-2x^2=3\]\[x+y=3\]

OpenStudy (anonymous):

from 2º equation y=3-x substituting this in the 1º equation: (3-x)^2-2x^2=3 9-6x+x^2-2x^2=3 -x^2-6x+9=3 x^2+6x-6=0 solve this using quadratic formula